OpenStudy (anonymous):

What are the domain and range of the function y=2√3x+4-5?

OpenStudy (ash2326):

Is this your question \[\large y= 2\sqrt{3x+4}-5\]

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Square root can't have a negative number inside so \[3x+4\ge 0\] or \[x\ge -\frac{4}{3}\] so Domain of y is \[[-\frac{4}{3}, \infty]\]

OpenStudy (ash2326):

minimum value of y will occur when x=-4/3, that's when the root will be zero \[y=0-5=-5\] so this is the minimum value of y, as x increases, y will increase so range of y is \[\large [-5, \infty)\]

OpenStudy (anonymous):

?

OpenStudy (ash2326):

Sorry the domain will have the round bracket to the right side \[[-\frac{4}{3}, \infty)\]

OpenStudy (anonymous):

it isn't clear where the square root is suppose to end, but 2*sqrt(3x+4)-5 seems more likely than 2*sqrt(3x+4-5) or 2*sqrt(3x)+4-5. The only requirement we have for x is that it makes the number in the radical >= 0: 3x+4 >= 0 3x >= -4 x >= -4/3 The range is the set of numbers that y could be. The smaller x is, the smaller y is, so the smallest y can be is when x is as small as it can be: x = -4/3 makes y = 2*0 - 5 = -5 y can be bigger if x gets bigger, so y >= -5

OpenStudy (anonymous):

now do you get it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome :)